1.Calculate the number of moles of avidin added for each titration based on the added volume of avidin stock solution and a MW of 68,000 g/mol.

Conduct the following steps for each titration:

1) Add Vbuffer of 0.02 M sodium phosphate buffer, pH 7, to a disposable plastic cuvette.

2) Add Vavidin of a solution of 10 mg/mL avidin in 0.02 M sodium phosphate buffer, pH 7. Mix thoroughly by micropipetting up and down.

3) Put the cuvette in the spectrophotometer and zero it at 500 nm.

4) Add VHABA,i of a 1 mM solution of HABA in 0.02 M sodium phosphate buffer, pH 7to the cuvette. Mix thoroughly by micropipetting up and down.

5) Measure the absorbance at 500 nm. 6) Repeat Steps 4 & 5 five additional times (i.e., until you have added a total of 24 µL of the HABA solution for Titration 1 and 42 µL of this solution for Titration 2).

7) Add VHABA,f of the HABA solution and mix thoroughly by micropipetting up and down. Record the absorbance as Asat.

8) Add Vbiotin of a 1 mM solution of biotin. Mix thoroughly by micropipetting up and down. Record the absorbance.


2. Describe how Describe how the equilibrium association constant can be determined from a Scatchard plot of the ratio of fractional saturation to equilibrium concentration of ligand vs. fractional saturation.
Answer #1

1. Concentration of Avidin = 10mg/mL

That is in 1mL solution there is 10mg Avidin.(10mg = 0.010g as 1mg = 0.001g)

Moles of Avidin per mL solution = mass(g)/molar mass(g/mol) = 0.010g/68000g/mol = 1.47×10^-7 moles per mL solution.

Moles of Avidin added in each titration = volume of solution added in mL × (1.47×10^-7 moles/mL) .

You have to just put the volume of Avidin solution added in each titration in above formula to get the exact moles of Avidin added.